A solenoid is 3 cm long and has a radius of 0.50 cm. The wire carries 2.0 A of current. The magnetic field at the center is A. 0.0419 T b. 0.099 T C. 0.0013 T D. 20 T


1
Expert's answer
2022-03-18T12:00:52-0400

The magnetic field inside solenoid

"B=\mu_0 \frac{n}{\sqrt{4r^2+l^2}}I"

"B=4\pi\times 10^{-7}\times \frac{1/0.03}{\sqrt{4\times 0.005^2+0.03^2}}\times2=0.0419\:\rm T"

Answer: A. 0.0419 T